# Extracting Participation in Collective Action from Social Media

### v1

@@ Introduction
The societal significance of online-mediated behavior change is growing, with grassroots, large-scale transformations being increasingly recognized as critical to addressing contemporary global challenges (Van Vugt 2009; Otto et al. 2020). While some online movements fail to yield tangible outcomes, in multiple occasions the Social Web has demonstrated its transformative potential in mobilizing collective action (Gerbaudo 2012) to achieve various political, societal, and economic goals (Yasseri et al. 2016; Chiovaro et al. 2021; Boulianne, Lalancette, and Ilkiw 2020; Lucchini et al. 2022).
@@
To address these challenges, we combined insights from literature on mobilization and activism with modern Natural Language Processing (NLP) techniques to develop open-source text classifiers to detect expressions of participation in collective action from social media posts [^1] Data and code are available at https://tinyurl.com/4tj57wfk. Models are shared at https://huggingface.co/ariannap22.. Our classifiers identify the presence of participation and categorize it into one of four levels: (i) identifying collective problems and proposing solutions, (ii) engaging in calls-to-action, (iii) expressing intent to act, and (iv) reporting actual involvement in action. The tools are topic-agnostic, and therefore can be applied to large-scale platforms to generate extensive and granular ground truth data for various applications.
@@
We found that small models can reliably quantify participation in collective action, often rivaling Large Language Models (LLMs). When applied to the binary classification task of identifying any expression of participation, the best BERT-based model achieved a macro F1 score of $0.71$, close to the $0.77$ of the best-performing LLM approach. In the multi-class classification of participation levels, the best BERT model obtained a macro F1 score of $0.44$, not far from the $0.52$ obtained by the best fine-tuned LLM. Notably, while fine-tuned LLMs excel at nuanced classifications, they demand significantly more computational resources, with inference times up to two orders of magnitude larger than what smaller transformer-based models require. The choice of model should therefore balance resource availability with task requirements.
@@
We began our analysis by considering the top 40,000 subreddits ranked by all-time subscriber count, from which we identified a subset of 73 subreddits that contained the terms “activism,” “activist,” or “rights” in their titles or descriptions. Following a manual review, we selected 42 subreddits pertinent to our study’s scope, containing a total of 120 million comments posted from the creation of these subreddits through December 2023 (see Section A.1 in Appendix A for the exclusion criteria and the list of selected subreddits).
@@
Overall, we employed five quality-control strategies. First, we presented the tasks as images rather than HTML text to prevent easy task automation. Second, we only recruited “master” workers who had completed at least 5,000 annotations on MTurk with a minimum 95% acceptance rate. Third, we discarded all annotations from workers failing more than 50% of control samples. Fourth, we removed samples with fewer than two annotators after filtering. Finally, we excluded samples without a clear majority vote. The final inter-annotator agreement measured with Krippendorff’s alpha is 0.86. The resulting annotated dataset consists of 369 samples: 100 labeled as None and 269 as expressing Participation in collective action, with 202 labeled as Problem-solution, 44 as Call-to-action, 14 as Intention, and 9 as Execution.
@@
To address data sparsity in the annotated data, particularly in classes representing higher levels of participation, we applied two data augmentation techniques (Table 1).
The first approach involved generating synthetic samples using Llama3 8B Instruct (Møller et al. 2024). Excluding the majority class (Problem-solution), we generated 20 synthetic examples for each comment in the Call-to-action, Intention, Execution, and None classes. This dataset is referred to as Syn A. We also created a variation focusing solely on the minority classes Intention and Execution (Syn I/E). All generated samples were manually reviewed for accuracy. The generation prompt is provided in Appendix C.
@@
Table 1: N. of participation labels by augmentation type.
@@
For binary classification, we used the dictionary of collective action words (Smith, McGarty, and Thomas 2018) to calculate the fraction of such words in the text. We then selected the threshold maximizing the difference between the True Positive Rate (TPR) and False Positive Rate (FPR) on the training set.
@@ Results
Table 2 reports the performance of the binary classification task–the first step in our layered approach–measured in terms of F1 scores. The best-performing models were the zero-shot Llama3 model (both with and without the collective action problem definition in the prompt) and the RoBERTa classifier trained on the synthetically augmented training set. Across methods, the synthetic augmentation improved the classification performance considerably, up to $80\%$ for the macro F1 for RoBERTa. Fine-tuned LLMs underperformed simpler approaches.
@@
Table 2: F1 scores for the binary classification task. The three best-performing models are in bold.
Method Train. set Part. None Macro Avg. Weight. Avg.
@@
Syn A (no def) 0.50 0.59 0.54 0.57
Syn A (def) 0.51 0.59 0.55 0.57
DPO Syn A (def) 0.47 0.42 0.44 0.43
Table 3: Performance metrics for the multi-class classification task as a second step of the layered approach. The three best-performing models are in bold.
@@
Table 3 reports performances for the multi-class task, considering the inclusion of the collective action problem definition in the LLM prompts. The SFT Llama3 model, trained on crowdworker-annotated data with synthetic augmentation for all classes except the majority class (CS+Syn A), achieved the best overall performance. The DPO Llama3 model, also trained on annotated data with the two synthetic augmentation strategies, followed closely. Given that prompt variations and the limited number of Intention and Execution labels in the test set affect reliability for these classes, the RoBERTa classifier, trained on synthetically augmented data, remains a robust choice for this stage. However, the SFT and DPO models consistently outperformed others, particularly in classifying the minority classes Intention and Execution.
The ranking of models remained largely consistent whether or not the collective action problem definition was included in the LLM prompts or when models were applied directly to the multi-class task without a layered approach (see Appendix B, Tables B2 and B3).
@@
Figure 1: Validation of the proposed approach. (a-d) comparison with topic modeling; (e) comparison with stance detection; (f) impact on classification performance given by the use of collective action keywords to define training and test sets.
@@
We used the collective action words dictionary (Smith, McGarty, and Thomas 2018) to filter comments in both the training and test sets, as well as to extract relevant portions of each comment for classification models and annotators. To ensure our classifier does not rely solely on the presence of these words, we tested the classifier on four variations of the test set: (i) removal of action dictionary words, (ii) substitution of action dictionary words with random tokens from the model vocabulary, (iii) removal of random words, and (iv) substitution of random words with random tokens.
@@
To demonstrate the broader applicability of our framework, we applied the classifier to detect participation in collective action in a subset of UK parliamentary debate transcripts (MySociety 2024). We selected debates from March 22nd, 25th, and 26th, 2024, totaling 1,000 text samples. Our approach effectively distinguishes messages expressing participation from those that do not, either by being unrelated to collective action or not identifying the issues as problematic (Table 4).
@@
Figure 2: Comparison of climate change-focused comment percentages and participation in collective action levels across subreddits, with x-axis values normalized using min-max scaling.
@@
We now demonstrate the effectiveness of our method for detecting comments indicative of participation in collective action, comparing it to widely used approaches in the literature. We used climate action as a case study due to the extensive research in this area, particularly within Reddit (Oswald and Bright 2022; Treen et al. 2022). Previous studies have largely relied on subreddit selection and keyword filtering, assuming that the presence of certain keywords or subreddit participation directly reflects participation in climate action (Parsa et al. 2022; Lenti et al. 2024). In contrast, we show that subreddits selected based on keyword counts may not always align with actual participation in collective action.
@@
With the rise of online media in public discourse (Boulianne and Theocharis 2020), research increasingly seeks to characterize the socio-political attributes of social media users through embedding-based methods that distill semantic information from user content to infer attributes like age, income, and partisanship (Aceves and Evans 2024). These embeddings have been used to study political engagement by correlating key dimensions with outcomes such as collective decision-making, online mobilization, and advocacy for global causes (Israeli, Kremiansky, and Tsur 2022; Kumar et al. 2018; Lenti et al. 2024). However, such studies often rely on coarse proxies, such as subreddit participation as a stand-in for activism on Reddit (Lenti et al. 2024).
@@
Drawing inspiration from the literature on collective action and mobilization, we contribute to the Computational Social Science community with new open-source text classifiers designed to accurately extract expressions of participation in collective action across topics. Our classification pipeline can operate in a layered fashion to capture different levels of participation. We achieve this by employing state-of-the-art Natural Language Processing models through a supervised learning approach, complemented by modern data augmentation techniques. We show that relatively high classification accuracy is attainable even with a limited number of manually-labeled examples. In our analysis of 61 different combinations of models and data augmentation strategies, we discovered that transformer-based small language models can compete effectively with more modern Large Language Models. Notably, our BERT-based classifier offers an optimal balance between performance, computational efficiency, and cost-effectiveness across both layers of classification (RQ1).
@@
Our proposed method has some limitations. First, the volume of the annotated training set is relatively small. Although we have implemented data augmentation techniques to address this limitation, the collection of additional annotations could further enhance classification quality. Second, while the training set consists of crowdsourced annotations, our test set is annotated by domain experts. Although we prioritized the accuracy and robustness of the test set, incorporating expert-annotated data into the training set could bolster its reliability. Third, while we framed the classification of nuanced levels of participation in collective action as a multi-class task, it could alternatively have been approached as a multi-label task, given that multiple levels may apply to a single text. Last, our experiments are almost entirely based on Reddit data, which calls for more extensive investigation of the generality of our approach across platforms and contexts.
@@
Related Work
In contrast, classifiers based on text embeddings have been used to extract social dimensions such as emotions, empathy (Choi et al. 2020, 2023), and specific stages of collective action like calls-to-action (Rogers, Kovaleva, and Rumshisky 2019). While these methods provide greater flexibility, their performance is often constrained by their generalizability across contexts.
The advent of Large Language Models (LLMs) has created new possibilities for computational social science, especially in data-scarce scenarios (Møller and Aiello 2024). While LLMs offer broad utility, their effectiveness relative to fine-tuned classifiers depends on the task and domain, with trade-offs in annotation and computational costs. In some cases, simpler fine-tuned classifiers outperform LLMs in specific tasks (Ziems et al. 2024), underscoring the importance of task-specific model selection.
Reddit has been a valuable source for analyzing group-specific characteristics, such as mental health (Kim et al. 2023) and political leanings (Hofmann, Schütze, and Pierrehumbert 2022). However, individual engagement in activism through textual traces remains underexplored. Studies often use proxies like subreddit involvement to measure activism (Lenti et al. 2024), overlooking comment-level commitments to collective action.
@@ A.2 Crowdsourcing Task
Annotation Task Instructions
Given the social media comment, which is the most appropriate label to represent the expressed level of participation in collective action? Refer to the ’Instructions’ panel.
Instructions
We are interested in collective action expressed in social media comments. Collective action can refer to a broad range of topics, such as environment, civil rights and mental health, and can reflect a spectrum of stances. Consider the given labels, roughly defined on a scale of increasing level of participation in collective action:
  • Problem-solution: the comment identifies an issue relevant to collective action as problematic, potentially assigning blame to the responsible and proposing a view on how to address the problem.
  • Call-to-action: the comment attempts to convince others to join or support a cause (e.g. a “call to arms”).
  • Intention: the comment expresses the author’s willingness or interest to act.
  • Execution: the comment directly reports the author’s involvement in some form of action.
  • None: None of the above.
Annotate the social media comment below according to the labels. To ease the labeling process, if multiple labels apply please select the highest on the scale (e.g. if both “Intention” and “Execution” apply, select “Execution”). Please press ’More Instructions’ and read an example for each label before annotating. Any suspected fraud or abuse will result in forfeiture of compensation.
Examples for Each Label
  • Problem-solution: “The decrease gender violence, one should start from school education and then offer continuous support through the workplace.”
  • Call-to-action: “The wealth gap is widening every day, and it’s time for action. Join us in calling for economic justice by attending the rally downtown this Saturday.”
  • Intention: “I wish to do my part in fighting racial injustice and join some community meetings.”
  • Execution: “Just got back from volunteering at the local food bank. Small actions can make a difference.”

### v2

@@ Introduction
The societal significance of online-mediated behavior change is growing, with grassroots, large-scale transformations being increasingly recognized as critical to addressing contemporary global challenges (Van Vugt 2009; Otto et al. 2020). While some online movements fail to yield tangible outcomes, in multiple occasions the Social Web has demonstrated its transformative potential in mobilizing collective action (Gerbaudo 2012) to achieve various political, societal, and economic goals (Yasseri et al. 2016; Boulianne and Theocharis 2020; Lucchini et al. 2022).
@@
To address these challenges, we combined insights from literature on mobilization and activism with modern Natural Language Processing (NLP) techniques to develop open-source text classifiers to detect expressions of participation in collective action from social media posts [^1] Data and code are available at https://github.com/ariannap13/extract˙collective˙action. Models are shared at https://huggingface.co/ariannap22.. Our classifiers identify the presence of participation and categorize it into one of four levels: (i) identifying collective problems and proposing solutions, (ii) engaging in calls-to-action, (iii) expressing intent to act, and (iv) reporting actual involvement in action. The tools are topic-agnostic, and therefore can be applied to large-scale platforms to generate extensive and granular ground truth data for various applications.
@@
We found that small models can reliably quantify participation in collective action, often rivaling Large Language Models (LLMs). When applied to the binary classification task of identifying any expression of participation, the best BERT-based model achieved a macro F1 score of $0.65$, close to the $0.71$ of the best-performing LLM approach. In the multi-class classification of participation levels, the best BERT model obtained a macro F1 score of $0.44$, not far from the $0.52$ obtained by the best fine-tuned LLM. Notably, while fine-tuned LLMs excel at nuanced classifications, they demand significantly more computational resources, with inference times up to two orders of magnitude larger than what smaller transformer-based models require. The choice of model should therefore balance resource availability with task requirements.
@@
Related Work
Social Movements and Collective Action
Language Models in Text Classification
Classifiers based on text embeddings have been used to extract social dimensions such as emotions, empathy (Choi et al. 2020, 2023), and specific stages of collective action like calls-to-action (Rogers, Kovaleva, and Rumshisky 2019). While these methods provide greater flexibility, their performance is often constrained by their generalizability across contexts and scopes.
The advent of Large Language Models (LLMs) has created new possibilities for computational social science, especially in data-scarce scenarios (Møller and Aiello 2024). Prior work has explored their utility in tasks such as stance detection (Lan et al. 2024), hate speech identification (Kumarage, Bhattacharjee, and Garland 2024), and misinformation detection (Huang et al. 2025), often stressing the importance of hybrid approaches in such tasks. However, less attention has been paid to the ability of LLMs to classify user participation in collective action, a task requiring a nuanced understanding of social movement dynamics. Unlike previous studies that focus on a single classification paradigm, our work systematically compares dictionary-based, centroid-based, BERT-based, and LLM-based classifiers to assess their effectiveness in capturing participation signals. This comparative approach allows us to evaluate generalizability across different modeling strategies, shedding light on the trade-offs between interpretability, accuracy, and computational cost in collective action classification. Specifically, while LLMs offer broad utility, their effectiveness relative to fine-tuned classifiers depends on the task and domain, with trade-offs in annotation and computational costs. In some cases, simpler fine-tuned classifiers outperform LLMs in specific tasks (Ziems et al. 2024), underscoring the importance of task-specific model selection.
Studying Activism on Reddit
Reddit has been a valuable source for analyzing group-specific characteristics, such as mental health (Kim et al. 2023) and political leanings (Hofmann, Schütze, and Pierrehumbert 2022). However, individual engagement in activism through textual traces remains underexplored. Existing studies often use proxies like subreddit involvement to measure activism (Lenti et al. 2025), overlooking comment-level commitments to collective action. In this work, we address this limitation by focusing on the extraction of comment-level participation in collective action, offering a new perspective on how online communities express engagement in activism.
@@
We began our analysis by considering the top 40,000 subreddits ranked by all-time subscriber count, from which we identified a subset of 73 subreddits that contained the terms “activism,” “activist,” or “rights” in their titles or descriptions. The identification of these subreddits was performed automatically by searching for the presence of these seed terms, which were selected for their direct relevance to the topic of collective action. Following a manual review, we selected 42 subreddits pertinent to our study’s scope, containing a total of 120 million comments posted from the creation of these subreddits through December 2023 (see Section A.1 in Appendix A for the exclusion criteria and the list of selected subreddits).
@@
Overall, we employed five quality-control strategies. First, we presented the tasks as images rather than HTML text to prevent easy task automation. Second, we only recruited “master” workers who had completed at least 5,000 annotations on MTurk with a minimum 95% acceptance rate. Third, we discarded all annotations from workers failing more than 50% of control samples. Fourth, we removed samples with fewer than two annotators after filtering. Finally, we excluded samples without a clear majority vote. The final inter-annotator agreement measured with Krippendorff’s alpha is 0.86. The resulting annotated dataset consists of 369 samples: 100 labeled as None and 269 as expressing Participation in collective action, with 202 labeled as Problem-solution, 44 as Call-to-action, 14 as Intention, and 9 as Execution. Later on, we will refer to this dataset as CS.
@@
To address data sparsity in the annotated data, particularly in classes representing higher levels of participation, we applied two data augmentation techniques.
The first approach involved generating synthetic samples using Llama3 8B Instruct (Møller et al. 2024). Excluding the majority class (Problem-solution), we generated 20 synthetic examples for each comment in the Call-to-action, Intention, Execution, and None classes. This dataset is referred to as Syn A. We also created a variation focusing solely on the minority classes Intention and Execution (Syn I/E). The generation prompt is provided in Appendix C. All generated samples were manually reviewed for accuracy using a pipeline based on the validity criteria from the prompt for each anchor example (e.g., semantic similarity, structure, meaning, intent, and key details). A table was created to binary-check the samples against these criteria.
@@
Table 1 summarizes the label volumes for the crowdsourced dataset (CS), the synthetic augmentation for minority classes (CS + Syn I/E), the synthetic augmentation for all classes (CS + Syn A), the Reddit-informed extension (Ext), and the combination of both Reddit extension and synthetic augmentation for minority classes (Ext + Syn I/E).
@@
For binary classification, we used the dictionary of collective action words (Smith, McGarty, and Thomas 2018) to calculate the fraction of such words in the text. We then selected the threshold maximizing the difference between the True Positive Rate (TPR) and False Positive Rate (FPR) on the training set, prioritizing a class-balanced decision boundary to avoid under-detecting instances of “Participation”. While we initially considered maximizing the F1-score, we found that it tended to favor the majority class, leading to a threshold that disproportionately benefited the more frequent “None” class. In contrast, maximizing $TPR-FPR$ allowed for a more balanced threshold, ensuring better detection of “Participation” cases without bias towards the majority class.
@@ Results
Table 2 reports the performance of the binary classification task–the first step in our layered approach–measured in terms of F1 scores. Details on precision and recall values can be found in Table  B2 in Appendix B. The best-performing models were the zero-shot Llama3 model (both with and without the collective action problem definition in the prompt) and the RoBERTa classifier trained on the synthetically augmented training set. Across methods, the synthetic augmentation improved the classification performance considerably, up to $80\%$ for the macro F1 for RoBERTa. Fine-tuned LLMs underperformed simpler approaches.
@@
Method Train. set Part. None Macro Avg. Weigh. Avg.
@@
CS + Syn A (no def) 0.50 0.59 0.54 0.57
CS + Syn A (def) 0.51 0.59 0.55 0.57
DPO CS + Syn A (def) 0.47 0.42 0.44 0.43
Table 2: F1 scores for the binary classification task, with the top three models in bold. CS denotes the crowdsourced dataset, CS + Syn A indicates synthetic augmentation for all classes, and def and no def refer to whether the characteristics of a collective action problem are defined in the LLM prompt or not.
@@
Table 3: F1 scores for the multi-class classification task as a second step of the layered approach, with the top three models in bold. CS denotes the crowdsourced dataset, CS + Syn I/E indicates synthetic augmentation for minority classes, CS + Syn A for all classes, Ext denotes the Reddit-informed extension, and Ext + Syn I/E defines the combination of both Reddit extension and synthetic augmentation for minority classes. All LLM prompts define a collective action problem.
@@
Table 3 reports performances for the multi-class task in terms of F1 scores, considering the inclusion of the collective action problem definition in the LLM prompts. Details on precision and recall values can be found in Table  B3 in Appendix B. The SFT Llama3 model, trained on crowdworker-annotated data with synthetic augmentation for all classes except the majority class (CS+Syn A), achieved the best overall performance. The DPO Llama3 model, also trained on annotated data with the two synthetic augmentation strategies, followed closely. Given that prompt variations and the limited number of Intention and Execution labels in the test set affect reliability for these classes, the RoBERTa classifier, trained on synthetically augmented data, remains a robust choice for this stage. However, the SFT and DPO models consistently outperformed others, particularly in classifying the minority classes Intention and Execution.
The ranking of models remained largely consistent whether or not the collective action problem definition was included in the LLM prompts or when models were applied directly to the multi-class task without a layered approach (see Appendix B, Tables B4 and B5).
@@
Figure 1: Validation of the proposed approach. (a-d) comparison with topic modeling; (e) comparison with stance detection – A for agree, D for disagree, N for neutral; (f) impact of collective action keywords on classification performance.
@@
We used the collective action words dictionary (Smith, McGarty, and Thomas 2018) to filter comments in both the training and test sets, as well as to extract relevant portions of each comment for classification models and annotators. To ensure our classifier does not rely solely on the presence of these words, we tested the classifier on four variations of the test set: (i) removal of action dictionary words, (ii) substitution of action dictionary words with random tokens from the model vocabulary, (iii) removal of random words, and (iv) substitution of random words with random tokens. This approach aligns with established practices in NLP for evaluating model robustness using adversarial datasets (Ebrahimi et al. 2017; Ghaddar et al. 2021).
@@
To demonstrate the broader applicability of our framework, we applied the classifier to detect participation in collective action in a subset of UK parliamentary debate transcripts (MySociety 2024). We selected debates from March 22nd, 25th, and 26th, 2024, totaling 1,000 text samples. Our approach effectively distinguishes messages expressing participation from those that do not, either by being unrelated to collective action or not identifying the issues as problematic (Table 4). While this analysis is preliminary, it demonstrates the potential of our approach for structured political discourse, highlighting a promising avenue for future research on collective action detection in institutional settings.
@@
Figure 2: Comparison of climate change-focused comment percentages and participation in collective action levels across subreddits, x-axis values normalized.
@@
We now demonstrate the effectiveness of our method for detecting comments indicative of participation in collective action, comparing it to widely used approaches in the literature. We used climate action as a case study due to the extensive research in this area, particularly within Reddit (Oswald and Bright 2022; Treen et al. 2022). Previous studies have largely relied on subreddit selection and keyword filtering, assuming that the presence of certain keywords or subreddit participation directly reflects participation in climate action (Parsa et al. 2022; Lenti et al. 2025). In contrast, we show that subreddits selected based on keyword counts may not always align with actual participation in collective action.
@@
With the rise of online media in public discourse (Boulianne and Theocharis 2020), research increasingly seeks to characterize the socio-political attributes of social media users through embedding-based methods that distill semantic information from user content to infer attributes like age, income, and partisanship (Aceves and Evans 2024). These embeddings have been used to study political engagement by correlating key dimensions with outcomes such as collective decision-making, online mobilization, and advocacy for global causes (Israeli, Kremiansky, and Tsur 2022; Kumar et al. 2018; Lenti et al. 2025). However, such studies often rely on coarse proxies, such as subreddit participation as a stand-in for activism on Reddit (Lenti et al. 2025).
@@
Drawing inspiration from the literature on collective action and mobilization, we contribute to the Computational Social Science community with new open-source text classifiers designed to accurately extract expressions of participation in collective action across topics. Our classification pipeline can operate in a layered fashion to capture different levels of participation. We achieve this by employing state-of-the-art Natural Language Processing models through a supervised learning approach, complemented by modern data augmentation techniques. We show that relatively high classification accuracy is attainable even with a limited number of manually-labeled examples. In our analysis of 61 different combinations of models and data augmentation strategies, we discovered that transformer-based small language models can compete effectively with more modern Large Language Models. Notably, our BERT-based classifier offers an optimal balance between performance, computational efficiency, and cost-effectiveness across both layers of classification (RQ1). While the specific reasons behind the comparable performance of small language models and fine-tuned LLMs on our binary classification task remain unclear, similar findings have been reported in both NLP and Computational Social Science literature (Bucher and Martini 2024; Ziems et al. 2024). This suggests that smaller, domain-specific models can rival larger models for certain tasks, and we consider this an important avenue for future research.
@@
Limitations and Future Work
Our proposed method has some limitations. First, our focus on a limited number of seed terms—“activism,” “activist,” and “rights”—may have biased the selection of subreddits, potentially overlooking relevant but ideologically diverse communities. This could skew the dataset toward left-leaning discourse, affecting downstream applications. Future work could refine subreddit selection by incorporating additional terms or applying more systematic, politically balanced criteria. Second, the volume of the annotated training set is relatively small. While we have validated the augmented data and used it strategically to mitigate the cost of crowdsourcing annotations, its influence on model performance and potential biases warrants further investigation. This is particularly important given class imbalance and the presence of rare classes, underscoring the need for future research on their treatment and the development of tailored augmentation strategies to ensure robustness and reliability. Third, while the training set consists of crowdsourced annotations, our test set is annotated by domain experts. Although we prioritized the accuracy and robustness of the test set, incorporating expert-annotated data into the training set could bolster its reliability. Fourth, while we framed the classification of nuanced levels of participation in collective action as a multi-class task, it could alternatively have been approached as a multi-label task, given that multiple levels may apply to a single text. Fifth, our experiments are primarily based on Reddit data, with the training set exclusively sourced from Reddit. While this provides a sufficient foundation for an initial study, further research is needed to explore the generalizability of our approach across different platforms and contexts. Finally, while we provide an extensive analysis of model selection, configuration design, and training set variations, our exploration is not exhaustive. Future research could expand on these aspects by experimenting with additional modeling choices and training strategies.
Potential Applications
@@
Acknowledgments
We acknowledge the support from the Carlsberg Foundation through the COCOONS project (CF21-0432). The funder had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@ A.2 Crowdsourcing Task
For annotation task instructions, see crowdsource_instructions.pdf in the GitHub repo https://github.com/ariannap13/extract˙collective˙action.